Solve the problem involving the third law of Kepler's planetary motion with complete solutions. Determine the period of a pendulum using the following lengths:

A) 4.0 m

B) 7.5 m

C) 15 m


1
Expert's answer
2021-12-22T14:12:39-0500

The period of oscillation is:


"T_1=2\pi\sqrt{l_1/g}=4\text{ s};\\\space\\
T_2=2\pi\sqrt{l_2/g}=5.5\text{ s};\\\space\\
T_3=2\pi\sqrt{l_3/g}=7.8\text{ s}."